# Progressive multi-fidelity learning for physical system predictions

### v1

@@ Abstract
Multi-fidelity , data fusion , progressive learning , surrogate model , scientific machine learning
@@ 1 Introduction
In these scenarios, multi-fidelity strategies emerge as a powerful solution in many areas of science and engineering [11, 12, 13, 14, 15]. By leveraging data of different fidelity levels, multi-fidelity approaches enhance the model’s ability to generalize across different regions, particularly where high-fidelity information is limited or unavailable, thereby achieving an effective compromise between computational cost and model accuracy. A wide range of multi-fidelity surrogate modeling techniques are developed using Gaussian processes [16, 17, 18, 19], and, more recently, deep learning methods [20, 21, 22, 23, 24], which address key challenges such as high dimensionality, scalability, and strongly nonlinear correlations across datasets. Nevertheless, most existing approaches are tailored to settings where datasets are structurally homogeneous – for example, solution data generated with the same solver or model at varying discretization levels, or data from different solvers producing the same output quantities of interest [12]. Therefore, challenges arise in incorporating diverse sources of information or different data types, a problem often referred to in machine learning as multi-modal learning, impacting a wealth of application fields ranging from cancer biomarker discovery [25], precision medicine [26] and neuroimaging [27] to wearable sensors [28] and meteorological monitoring [29], just to mention a few. Additionally, datasets may not be available simultaneously at the time of surrogate construction but acquired at different stages, in which case retraining neural networks with new inputs may lead to the issue known as catastrophic forgetting.
To tackle these challenges, we propose a progressive multi-fidelity surrogate modeling paradigm, designed to incorporate new datasets of varying fidelity as they become available. This allows to progressively enhance prediction accuracy and reduce uncertainty, while effectively ensuring knowledge retention across updates. This model can sequentially incorporate diverse data types using tailored encoder neural networks. The encoders process and transform data into a meaningful latent representation that is suitable to be merged with the other fidelity datasets. The architecture of each encoder is selected according to the data modality: for example, feedforward networks for vector-valued data, recurrent networks (e.g., LSTMs [30]) for time series, convolutional networks [31] or vision transformers [32] for image data, and so on. The encoded features are concatenated and mapped to the target quantities of interest through a decoder, which is trained in a supervised manner on high-fidelity data.
@@
Moreover, this progressive multi-fidelity structure is beneficial not only in training but also in online deployment, as it adapts predictions to the inputs available at runtime, ensuring operational continuity. When higher-fidelity input data are unavailable due to time, budget, or sensor constraints, the model can seamlessly fall back to lower-fidelity levels. By leveraging the inductive bias of the progressive structure and the information contained in physically meaningful low-fidelity data, the framework mitigates the lack of physical consistency typical of data-driven methods in data-scarce regimes, thereby enhancing accuracy in generalization across time and parameters and improving reliability.
To evaluate the framework, we consider three applications spanning both (i) synthetic data obtained through the solution of nonlinear, time-dependent partial differential equations (PDEs) and (ii) real-world data. First, a reaction-diffusion system reconstructs high-fidelity spiral wave dynamics from coarse, noisy low-fidelity simulations. Second, a Navier–Stokes benchmark reconstructs vortex shedding from hierarchical low-fidelity inputs, including drag/lift coefficients, outlet sensors, and partial-domain snapshots. Third, an air pollution scenario estimates benzene concentrations from low-cost sensor data despite missing or unreliable inputs. These cases highlight key challenges: integrating multi-modal data, mitigating low-fidelity imperfections, and enabling reliable temporal and parametric extrapolation.
@@
  • Extensive validation across a range of PDE benchmarks with diverse quantities of interest, as well as a real-world air pollution case study, demonstrating the great flexibility of the proposed architecture and practical relevance in cases where noisy data are acquired from real sensor measurements.
@@ 2.1 Offline training
    ${\hat{\mathbf{y}}}^{(l)}={\hat{\mathbf{y}}}^{(l-1)}+\Psi^{(l)}({\mathbf{h}}_{\text{tot}}^{(l)})$ (3)
@@
The three steps outlined above define how input data are processed, at each level, to produce a set of progressively refined estimates $\{{\hat{\mathbf{y}}}^{(l)}\}_{l=0:L}$, approximating the high-fidelity target quantities of interest $\mathbf{y}$. At the $l$-th level, the optimal parameters ${\mathbf{W}}_{\Phi}^{(j)}\text{ and }{\mathbf{W}}_{\Psi}^{(j)}$ of the neural networks for encoding and decoding, $\Phi^{(l)}$ and $\Psi^{(l)}$ respectively, are obtained by solving the following optimization problem through back-propagation with ADAM [35] algorithm:
@@
This loss function is defined as the mean squared error (MSE) between the predicted output at level $l$ and the high-fidelity data, augmented with an additional regularization term denoted as $\lambda_{\text{reg}}$, intended to mitigate overfitting phenomena. Specifically, $L_{2}$ regularization is employed, i.e., the regularization term can be expressed as
$\lambda_{\text{reg}}\left({\mathbf{W}}_{\Phi}^{(j)},{\mathbf{W}}_{\Psi}^{(j)}\right)={\lambda_{\Phi}}\norm{{\mathbf{W}}_{\Phi}^{(j)}}_{2}^{2}+{\lambda_{\Psi}}\norm{{\mathbf{W}}_{\Psi}^{(j)}}_{2}^{2},\qquad\ \ \mbox{with}\ \lambda_{\Phi},\lambda_{\Psi}\in\mathbb{R}^{+}.$ (5)
@@
The preceding levels convey information to the subsequent ones through both the concatenation of latent variables and the additive feedback from the previous output. Consequently, each new level configures itself as a correction to be applied to the previous level, which, in contrast, remains unchanged.
@@ 2.2 Online prediction
Once the multi-fidelity networks are trained offline, they can be used to estimate high-fidelity outputs from the available low-fidelity inputs during the online prediction phase. The availability of low-fidelity input data depends on the online cost of assimilating or computing those data, given a certain computational budget or a time frame in which the progressive multi-fidelity model needs to operate. Therefore, if at testing time the budget allows for acquiring low-fidelity inputs $\{\mathbf{x}^{(l)}\}_{l\leq\bar{l}}$ up to level $\bar{l}$, the model will produce the corresponding prediction $\hat{\mathbf{y}}^{(\bar{l})}$. The online deployment procedure is summarized in Algorithm 1. Note that the level $\bar{l}$ may vary over time, allowing the model to navigate the hierarchy of outputs and ensure continuity in the prediction of the corresponding high-fidelity quantities.
The hierarchical ordering here is based on the availability of input data rather than on data accuracy, as in traditional multi-level or multi-grid methods, where all data come from a single source and the hierarchy is determined by a control parameter balancing accuracy and computational cost. In our framework, data may come from multiple sources or modalities, and the hierarchy is defined solely by the availability of the data at the prediction stage [40].
@@ 2.3 Case studies
  • Navier–Stokes PDE problem. A computational fluid dynamics benchmark leveraging hierarchical low-fidelity inputs (drag and lift coefficients, outlet sensors, and partial-domain snapshots) to reconstruct unsteady flow behavior.
@@ 3.1 Multi-fidelity setting
In this application, the quality of the input information determines the hierarchical order. Note that each level processes a dataset of different modality (vector valued, time-series, spatio-temporal snapshot data) through encoders tailored to the input data fusing their latent representation.
@@
    LF data are additionally contaminated by lognormal multiplicative noise with distribution $\mathcal{LN}{(0,0.8)}$ to simulate observation error.
To train offline the model, a small amount of expensive HF solution data are computed over a shorter time horizon $T_{\text{train}}=40<T=80$, at a limited set of parameter locations. In particular, $N_{\mu}=10$ $\mu$-values are selected over an equispaced grid of $\mathcal{P}$. We test the progressive MF method accuracy in predicting the HF solutions over an unseen set of parameters in $\mathcal{P}$ over the whole time window $[0,T]$. The input dimensions of the data at different fidelity levels are $d_{\text{in}}^{(0)}=2$, $d_{\text{in}}^{(1)}=4$, and $d_{\text{in}}^{(2)}=1024$, corresponding to the time and parameter values at Level 0, four vertex sensors at Level 1, and the low-fidelity spatial degrees of freedom at Level 2. The output dimension of the high-fidelity solution is $d_{\text{out}}=10000$. The dimensionality of both the Level 2 input and the high-fidelity output data is reduced using POD, retaining 9 and 11 modes, respectively, capturing about 90% of the total variance. Uncertainty bounds are computed from the statistical moments of trajectories obtained using ensemble-based methods over $m=30$ retrainings, each with randomly initialized network weights. More details about the implementation are provided in A.
@@ 3.2 Results
At level 1, temporal evolution data of the solution at the four vertices of the spatial domain are provided as additional inputs. Although this information is limited to the domain boundaries, we observe that incorporating physical insight leads to a significant improvement in prediction accuracy.
@@
Table 1: Relative error with respect to the HF reference solution, $\mathbf{y}$, is evaluated for the MF estimates, $\hat{\mathbf{y}}^{(l)}$, at each level $l$, as $\text{err}_{\%}^{(l)}=\frac{100\%}{N_{\text{test}}}\sum_{n=1}^{N_{\text{test}}}\frac{\norm{\mathbf{y}(t_{n},\mu_{n})-\hat{\mathbf{y}}^{(l)}(t_{n},\mu_{n})}_{2}}{\norm{{\mathbf{y}}(t_{n},\mu_{n})}_{2}}$, where $N_{\text{test}}$ is the number of time-parameter combinations in the test set.
@@
We observe that predictions within the training time interval $0,T_{\text{train}}]$, with $T_{\text{train}}=40$, are accurate at all levels, highlighting the neural network’s effectiveness as a surrogate for approximating high-dimensional, nonlinear systems in interpolation regimes – i.e., scenarios similar to those encountered during training. However, as we extrapolate beyond the training window into $T_{\text{train}},T]=[40,80]$, the predictive performance at level 0 deteriorates significantly. This behavior is expected, given that this level relies solely on the control parameters as input, which do not convey any dynamical or physical information about the system. Nevertheless, the model at level 0 reflects this limitation by providing wide uncertainty bounds, reflecting the increased epistemic uncertainty in this regime. In contrast, levels 1 and 2 show a marked improvement in forecasting performance, accurately tracking the system dynamics over long-term horizons for unseen parameters. This clearly demonstrates how the integration of physically meaningful LF data helps mitigate the lack of physical consistency often observed in purely data-driven models, thus enabling robust extrapolation capabilities. As expected, transitioning from level 1 to level 2 results in further gains in predictive accuracy and a noticeable reduction in uncertainty. This improvement, however, comes at the cost of running LF solvers over the full spatial domain.

### v2

# Progressive multi-fidelity learning with neural networks for physical system predictions

@@ Abstract
Multi-fidelity , data fusion , progressive learning , surrogate model , scientific machine learning , neural networks
@@ 1 Introduction
In these scenarios, multi-fidelity strategies emerge as a powerful solution in many areas of science and engineering [42, 7, 1, 24, 15]. By leveraging data of different fidelity levels, multi-fidelity approaches enhance the model’s ability to generalize across different regions, particularly where high-fidelity information is limited or unavailable, thereby achieving an effective compromise between computational cost and model accuracy. A wide range of multi-fidelity regression methods have been developed using Gaussian processes [40, 43, 3, 27], as well as other classical surrogate modeling techniques, including polynomial chaos expansions [38, 41]. More recently, deep learning has been proposed to address challenges related to high dimensionality, scalability, and strongly nonlinear cross-fidelity correlations [34, 31, 8, 18, 35]. Within this line of work, several learning strategies have been explored for multi-fidelity data fusion, including heterogeneous domain training [49], manifold alignment [44], diffusion models [51], and reinforcement learning approaches [9, 53]. Nevertheless, many existing approaches are tailored to settings in which datasets are structurally homogeneous. In this context, different fidelity levels provide information about the same quantity of interest, obtained from models or estimators of varying accuracy or computational cost. Typical examples include solution data generated using the same numerical solver at different resolution levels, or data produced by different solvers targeting the same output fields [7]. Therefore, challenges arise in incorporating diverse sources of information or different data types, a problem often referred to in machine learning as multi-modal learning, impacting a wealth of application fields ranging from cancer biomarker discovery [52], precision medicine [6] and neuroimaging [58] to wearable sensors [36] and meteorological monitoring [26], just to mention a few. Additionally, datasets may not be available simultaneously at the time of surrogate construction but acquired at different stages, in which case retraining neural networks with new inputs may lead to the issue known as catastrophic forgetting.
To tackle these challenges, we propose a progressive multi-fidelity surrogate modeling paradigm that incorporates new datasets of varying fidelity as they become available. This allows for progressively enhancing prediction accuracy and reducing uncertainty, while effectively ensuring knowledge retention across updates. This model can sequentially incorporate diverse data types using tailored encoder neural networks. The encoders process and transform data into a meaningful latent representation that is suitable to be merged with the other fidelity datasets. The architecture of each encoder is selected according to the data modality: for example, feedforward networks for vector-valued data, recurrent networks (e.g., LSTMs [17]) for time series, convolutional networks [28] or vision transformers [12] for image data, and so on. The encoded features are concatenated and mapped to the target quantities of interest through a decoder, which is trained in a supervised manner on high-fidelity data.
@@
Moreover, this progressive multi-fidelity structure is beneficial not only during training but also in online deployment. Since each fidelity level contributes an additive correction, the model can operate with only the input modalities available at inference time, by construction. When higher-fidelity inputs are unavailable due to time, budget, or data acquisition constraints, the framework systematically falls back to lower-fidelity levels. This enables automatic adaptation to asynchronous or missing inputs, ensuring operational continuity and progressively refined predictions as additional data become available. By leveraging the inductive bias of the progressive structure and the information contained in physically meaningful low-fidelity data, the framework mitigates the lack of physical consistency typical of data-driven methods in data-scarce regimes, thereby enhancing accuracy in generalization across time and parameters and improving reliability.
To evaluate the framework, we consider three applications spanning both (i) synthetic data obtained through the solution of nonlinear, time-dependent partial differential equations (PDEs) and (ii) real-world data. First, a reaction-diffusion system reconstructs high-fidelity spiral wave dynamics from coarse, noisy low-fidelity simulations. Second, a Navier-Stokes benchmark reconstructs vortex shedding from hierarchical low-fidelity inputs, including drag/lift coefficients, outlet sensors, and partial-domain snapshots. Third, an air pollution scenario estimates benzene concentrations from low-cost sensor data despite missing or unreliable inputs. These cases highlight key challenges: integrating multi-modal data, mitigating low-fidelity imperfections, and enabling reliable temporal and parametric extrapolation.
@@
  • Extensive validation across a range of PDE benchmarks with diverse quantities of interest, as well as a real-world air pollution case study, demonstrating the flexibility of the proposed architecture and practical relevance in cases where noisy data are acquired from real sensor measurements.
@@ 2.1 Offline training
    ${\hat{\mathbf{y}}}^{(l)}={\hat{\mathbf{y}}}^{(l-1)}+\Psi^{(l)}({\mathbf{h}}_{\text{tot}}^{(l)}){.}$ (3)
@@
The three steps outlined above define how input data are processed, at each level, to produce a set of progressively refined estimates $\{{\hat{\mathbf{y}}}^{(l)}\}_{l=0:L}$, approximating the high-fidelity target quantities of interest $\mathbf{y}$. At the $l$-th level, the optimal parameters ${\mathbf{W}}_{\Phi}^{(j)}\text{ and }{\mathbf{W}}_{\Psi}^{(j)}$ of the neural networks for encoding and decoding, $\Phi^{(l)}$ and $\Psi^{(l)}$ respectively, are obtained by solving the following optimization problem through back-propagation with the ADAM [25] algorithm:
@@
This loss function is defined as the mean squared error (MSE) between the predicted output at level $l$ and the high-fidelity data, augmented with an additional regularization term $\lambda_{\text{reg}}$ to mitigate overfitting, which is particularly important given the limited amount of high-fidelity training data. In this work, $L_{2}$ regularization is employed (rather than alternatives such as $L_{1}$ or dropout), as it provides a smooth penalty on the network weights and is well suited for regression tasks involving real-valued quantities. In particular, the regularization term is defined as
$\lambda_{\text{reg}}\left({\mathbf{W}}_{\Phi}^{(j)},{\mathbf{W}}_{\Psi}^{(j)}\right)={\lambda_{\Phi}}\norm{{\mathbf{W}}_{\Phi}^{(j)}}_{2}^{2}+{\lambda_{\Psi}}\norm{{\mathbf{W}}_{\Psi}^{(j)}}_{2}^{2},\qquad\lambda_{\Phi},\lambda_{\Psi}\in\mathbb{R}^{+}.$ (5)
@@
The preceding levels convey information to the subsequent ones through both the concatenation of latent variables and the additive feedback from the previous output. Consequently, each new level configures itself as a correction to be applied to the previous level, which, in contrast, remains unchanged. This additive formulation, together with freezing all previously trained components, enforces a one-directional flow of information, preserving earlier representations while higher-fidelity data contribute only via a learnable correction. In this sense, latent concatenation implements early fusion by integrating multi-fidelity inputs into a shared latent space, whereas additive corrections act as late fusion, progressively refining the output. This combination is expected to provide the benefits of both fusion strategies in the data integration process.
@@ 2.2 Online prediction
Once the multi-fidelity networks are trained offline, they can be used to estimate high-fidelity outputs from the available low-fidelity inputs during the online prediction phase. The availability of low-fidelity input data depends on the online cost of assimilating or computing those data, given a certain computational budget or a time frame in which the progressive multi-fidelity model needs to operate. Therefore, if at testing time the budget allows for acquiring low-fidelity inputs $\{\mathbf{x}^{(l)}\}_{l\leq\bar{l}}$ up to level $\bar{l}$, the model will produce the corresponding prediction $\hat{\mathbf{y}}^{(\bar{l})}$. The online deployment procedure is summarized in Algorithm 1. In practical time-series applications, some input modalities may not be consistently available at every time step, therefore the number of available inputs $\bar{l}$ can vary over time. Algorithm 1 is applied independently at each time step using the currently available modalities, producing a prediction $\hat{\mathbf{y}}^{(\bar{l})}$ that is progressively refined according to the inputs available at the given instant. This approach enables automatic adaptation to missing modalities, efficiently navigates the hierarchy of available data, and ensures continuity of operation.
The hierarchical ordering here is based on the availability of input data rather than on data accuracy, as in traditional multi-level or multi-grid methods, where all data come from a single source and the hierarchy is determined by a control parameter balancing accuracy and computational cost. In our framework, data may come from multiple sources or modalities, and the hierarchy is defined solely by the availability of the data at the prediction stage [30]. We note that the proposed online deployment assumes a strict hierarchy in data availability, i.e., low-fidelity inputs must be available consecutively from level 0 up to a maximum level $\bar{l}$. If data at a given level are missing, higher-level inputs cannot be exploited within the current progressive formulation, as the proposed strategy relies on the sequential concatenation of latent representations. Extensions aimed at relaxing this hierarchical constraint are briefly discussed in the final remarks in Section 6.
@@ 2.3 Case studies
  • Navier-Stokes PDE problem. A computational fluid dynamics benchmark leveraging hierarchical low-fidelity inputs (drag and lift coefficients, outlet sensors, and partial-domain snapshots) to reconstruct unsteady flow behavior.
@@ 3.1 Multi-fidelity setting
In this application, the quality of the input information determines the hierarchical order. Note that each level processes a dataset of different modality (vector valued, time-series, spatio-temporal snapshot data) through encoders tailored to the input data fusing their latent representation. We note that Level 0 corresponds to a standard single-fidelity surrogate setting, where the HF solution is learned using only HF data and the associated control parameters. The flexibility of the proposed framework allows the encoder and decoder components to be replaced, so that existing deep-learning-based HF surrogates (e.g., [13, 20]) can be embedded as the starting level in the hierarchy.
@@
    LF data are additionally contaminated by multiplicative log-normal noise to simulate observation error. In this work, we consider $\varepsilon\sim\mathrm{LogNormal}(0,0.8^{2}),$ i.e., $\varepsilon=\exp(\xi),\xi\sim\mathcal{N}(0,0.8^{2}).$
To train offline the model, a small amount of expensive HF solution data are computed over a shorter time horizon $T_{\text{train}}=40<T=80$, at a limited set of parameter locations. In particular, $N_{\mu}=10$ $\mu$-values are selected over an equispaced grid of $\mathcal{P}$. We test the progressive MF method accuracy in predicting the HF solutions over an unseen set of parameters in $\mathcal{P}$ over the whole time window $[0,T]$. The input dimensions of the data at different fidelity levels are $d_{\text{in}}^{(0)}=2$, $d_{\text{in}}^{(1)}=4$, and $d_{\text{in}}^{(2)}=1024$, corresponding to the time and parameter values at Level 0, four vertex sensors at Level 1, and the low-fidelity spatial degrees of freedom at Level 2. The output dimension of the high-fidelity solution is $d_{\text{out}}=10000$. The dimensionality of both the Level 2 input and the high-fidelity output data is reduced using POD, retaining 9 and 11 modes, respectively, capturing about 90% of the total variance. The POD bases are constructed from the same snapshot data used for training the surrogate model. Uncertainty bounds are computed from the statistical moments of trajectories obtained using ensemble-based methods over $m=30$ retrainings, each with randomly initialized network weights. More details about the implementation are provided in A.
@@ 3.2 Results
At level 1, temporal evolution data of the solution at the four vertices of the spatial domain are provided as additional inputs. Although this information is limited to the domain boundaries, it leads to a substantial improvement in prediction accuracy. In this setting, oscillations at the boundary are highly informative of the solution behavior over the entire domain, thereby explaining the marked performance gain obtained when boundary data are incorporated within the multi-fidelity framework.
@@
Table 1: Relative error with respect to the HF reference solution, $\mathbf{y}$, is evaluated for the MF estimates, $\hat{\mathbf{y}}^{(l)}$, at each level $l$, as $\text{err}_{\%}^{(l)}=\frac{100\%}{N_{\text{test}}}\sum_{n=1}^{N_{\text{test}}}\frac{\norm{\mathbf{y}(t_{n},\mu_{n})-\hat{\mathbf{y}}^{(l)}(t_{n},\mu_{n})}_{2}}{\norm{{\mathbf{y}}(t_{n},\mu_{n})}_{2}}$, where $N_{\text{test}}$ is the number of time-parameter combinations in the test set. Additionally, we report the relative error on the predicted POD expansion coefficients, shown in Fig. 4. This metric quantifies the surrogate model error in the reduced space and excludes the contribution of the spatial reconstruction error induced by POD truncation.
@@
We observe that, within the training time interval $0,T_{\text{train}}]$, with $T_{\text{train}}=40$, the predictions at the higher fidelity levels closely match the reference solution, indicating that the proposed architecture is effective in interpolation regimes. At level 0, noticeable deviations from the true response already arise within the training window, reflecting the intrinsic limitations of this level, which relies solely on control parameters and does not incorporate any dynamical or physical information. As the prediction horizon extends beyond the training interval into $T_{\text{train}},T]=[40,80]$, the performance at level 0 further deteriorates, as expected. This behavior is consistently accompanied by wide uncertainty bounds, which appropriately reflect the increased epistemic uncertainty associated with this lowest-fidelity model. In contrast, levels 1 and 2 show a marked improvement in forecasting performance, accurately tracking the system dynamics over long-term horizons for unseen parameters. This clearly demonstrates how the integration of physically meaningful LF data helps mitigate the lack of physical consistency often observed in purely data-driven models, thus enabling robust extrapolation capabilities. As expected, transitioning from level 1 to level 2 results in further gains in predictive accuracy and a noticeable reduction in uncertainty. This improvement, however, comes at the cost of running LF solvers over the full spatial domain.
The uncertainty estimates further enhance interpretation by highlighting regions where predictions are less reliable. Elevated uncertainty consistently aligns with regimes characterized by limited informative inputs (lower-fidelity information) or temporal extrapolation beyond the training window. In particular, at level 0 the uncertainty rapidly increases during long-term forecasting and spans the full oscillatory range of the POD coefficients, reflecting the model’s intrinsic limitations. At higher fidelity levels, uncertainty remains small in interpolation regimes and increases with the prediction horizon, capturing the accumulation of epistemic uncertainty in the recurrent LSTM forecasts. Overall, this behavior shows how the ensemble-based UQ acts as a practical safeguard against overconfident predictions in data-scarce or extrapolative settings.
We note that the POD basis used to reduce the dimensionality of the HF outputs imposes an intrinsic upper bound on the achievable accuracy of the reconstructed solution in physical space. To disentangle the different error contributions, we additionally report in Table 1 the relative error on the predicted POD expansion coefficients, which isolates the surrogate modeling error in reduced space and excludes the contribution of POD truncation. The observed reduction of POD coefficient error across fidelity levels mirrors the improvement in the reconstructed solutions, particularly at Levels 0 and 1, indicating that the dominant source of error in this example arises from the surrogate’s ability to predict the reduced dynamics rather than from POD truncation. At Level 2, the discrepancy between the coefficient error and the reconstructed solution error becomes non-negligible, reflecting the contribution of POD truncation. This effect could be mitigated by retaining a larger number of POD modes; however, this would require predicting a higher-dimensional output, typically associated with lower-variance modes exhibiting more complex temporal behavior, thereby increasing the difficulty of the regression task. In practice, we observed that increasing the number of retained modes from 11 to 16 led to a higher overall reconstruction error at level 2 (26.2% versus 18.6%), indicating that the chosen truncation represents a well balanced trade-off between expressiveness and learnability.